OpenStudy (anonymous):

(4 multipled by sq. root of 5) (3 multiplied by sq. root of 2) multiply, and simplify where possible. please show step by step

OpenStudy (lalaly):

\[(4\sqrt5)(3\sqrt2) = 4 \times 3 \times \sqrt{5 \times 2} = 12\sqrt10\]